# First moments of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions and their applications to subconvexity

### v1

@@ 2.8. Bessel functions
see, for example, [13, Proposition 8]. For any $x\in\mathbb{R}$, one also might write
@@
for any $j\geq 0$; see, for example, [45, Sections 16.12, 17.5 & 16.3]. Now, for any $h$, a smooth bump function supported on $[1/2,5/2]$ with bounded derivatives, we define
$\widehat{h}(t)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}^{+}}\frac{h(\sqrt{u})}{\sqrt{u}}\,e^{itu}\mathrm{d}u.$
We thus have (see [22, Corollary 8.2])
@@
$\sum_{k\equiv 0\bmod 2}i^{-k}h{\left(\frac{k-1}{K}\right)}J_{k-1}(x)=-\frac{K}{2\sqrt{x}}\,\text{Im}{\left(e{\left(\frac{x}{2\pi}-\frac{1}{8}\right)}\right)}\widehat{h}{\left(\frac{K^{2}}{2x}\right)}+O{\left(\frac{x}{K^{4}}\right)},$ (2.23)
@@
This shows that the oscillatory range for $J_{k-1}(x)$ is roughly $x>k^{2}$, when averaging over the weight $k$; within this range $J_{k-1}(x)$ obeys the asymptotic behavior as shown in (2.24). The function $\widehat{h}$ behaves like a Schwarz function, and satisfies that $\widehat{h}^{(j)}(x)\ll(1+|x|)^{-A}$ for any large $A$ and $j>0$ by repeated integration by parts sufficiently many times. On the other hand, if there does not exist the a factor $i^{k}$ in (2.24), one, however, has (see [22, Proposition 8.1])
@@
$\begin{split}4\sum_{k\equiv\iota\bmod 4}h{\left(\frac{k}{K}\right)}J_{k-1}(x)=&h{\left(\frac{k-1}{K}\right)}-i^{-\iota}\frac{K}{\sqrt{x}}\,\text{Im}{\left(e{\left(\frac{x}{2\pi}-\frac{1}{8}\right)}\right)}\widehat{h}{\left(\frac{K^{2}}{2x}\right)}\\ &+O{\left(\frac{x}{K^{3}}\left|h^{(3)}{\left(\frac{x}{K}\right)}\right|\right)}+O{\left(\frac{x}{K^{4}}\right)},\end{split}$ (2.24)
@@
Let $q$ be a prime and $M$ is also, satisfying that $(q,M)=1$. For any $f\in\mathcal{B}^{\ast}_{k}(M)$ and $F$ being a normalized $\text{GL}(3)$ Hecke-Maaß form of level $q^{2}$ with trivial nebentypus, we will consider the first moment of the central value $L(1/2,F\otimes f)$ in the levels aspects. Recall that $\mathcal{Q}_{F,f}\asymp M^{3}q^{4}$. The functional equation method shows that
$\begin{split}L(1/2,F\otimes f)=&\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}U{\left(\frac{n}{q^{2}M^{3/2}}\right)}\\ &\hskip 65.44142pt+\varepsilon(F\otimes f)\sum_{n\geq 1}\frac{A_{\widetilde{F}}(n,1)\lambda_{f}(n)}{\sqrt{n}}V{\left(\frac{n}{q^{2}M^{3/2}}\right)},\end{split}$
@@
$L\geq\frac{q^{2+\varepsilon}}{{\sqrt{M}}},\quad L\geq\frac{\sqrt{M}}{{q^{2-\varepsilon}}}$ (3.26)
for any $\varepsilon>0$. We now define
=$\mathscr{L}=\{p:p\text{ is prime, }L\leq p\leq 2L,(p,qM)=1\}.$
By equipping with the amplifier
$\mathfrak{A}_{f}=\sum_{\ell\in\mathscr{L}}\alpha_{\ell}\lambda_{f}(\ell),\quad\text{with }\alpha_{\ell}=\begin{cases}\lambda_{f}(\ell),&\text{ if }\ell\in\mathscr{L},\\ -1,&\text{ if }\ell=p^{2},\,p\in\mathscr{L},\\ 0,&\text{ otherwise},\end{cases}$ (3.27)
one attempts to estimate an amplified first moment
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}|\mathfrak{A}_{f}|^{2}L(1/2,F\otimes f).$
It is clear that, by the relation $\lambda^{2}_{f}(p)-\lambda_{f}(p^{2})=1$, one sees that $|\mathfrak{A}_{f}|\gg L/\log L$. If one opens the square and applies multiplicativity of Hecke eigenvalues, finally our main focus is on
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f).$ (3.28)
Note that here and thereafter, $\ell$ is either a prime or a square of a prime satisfying the coprimalities with the parameters $q$ and $M$.
@@
with $\mathcal{Q}^{1/2-\delta}_{F,f}\leq\mathcal{X}\leq\mathcal{Q}^{1/2+\varepsilon}_{F,f}$ for any $\delta,\varepsilon>0$. Here, $U$ is a smooth bump function supported $[1/2,5/2]$, identical to one for $x\in[1,2]$, with derivatives satisfying that $x^{j}U^{(j)}\left({x}/{\mathcal{X}}\right)\ll_{j}1$ for any $j\geq 0$.
By the Cauchy-Schwarz inequality, together with the Rankin-Selberg bounds for the second moments of Fourier coefficients, the trivial estimate for the double sum above is $\ll_{\varepsilon}\mathcal{X}^{1/2+\varepsilon}$. This is far away from the bound $\ll\mathcal{X}^{1/2+\varepsilon}/M$ leading to the convexity. The task is thus reduced to estimating (3.29) further. Our strategy next is the Petersson trace formula. To set up for our application of this, we trivially write the expression (3.29) above as
@@
It is straightforward to see that, by the fact $|\lambda_{f}(M)|\ll M^{-1/2}$ that the second term is bounded by $\ll_{\varepsilon}\sqrt{\mathcal{X}}/M^{3/2}$. For $(n\ell,M)=1$, notice that, by Lemma 2.6 one has the following relation
@@
Noticing that we are taking $k$ sufficiently large, we see that the sum over $c$ can be truncated at $c\leq\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$; otherwise, the contribution in the complementary range is negligibly small (i.e., $O_{A}(\mathcal{X}^{-A})$ for any sufficiently large $A$), with the help of the bound (2.20). By using the dyadic subdivision, we further consider the multiple sum
@@
with ${U}_{\ell}\left(y\right)=U(y)J_{k-1}\left({4\pi\sqrt{y\mathcal{X}\ell}}/{CM}\right)$ for any $y\in\mathbb{R}^{+}$. According to the properties of $J$-Bessel functions in §2.1, it follows that
@@
Next, we prepare to apply the Voronoĭ summation formula, Lemma 2.1. Before that, one needs to detect the coprimality between $\ell$ and the modulus of the twist $cM$. Recall $\ell$ is a prime or a square of a prime. To begin with, when $\ell$ is prime, the coprimality $(c,\ell)=1$ is ascertained by the first constraint of (3.26), which forces that $c<\ell$. Now, we turn to considering that the case of $\ell$ being a square of a prime. Observe that in the (possibly occurred) case where $\ell=p^{2},c=pc^{\prime}$, with $p\asymp L$, by (2.20), the quantity $\mathcal{S}(M,q)$ is negligible small, unless that $\sqrt{\mathcal{X}}\gg M^{1-\varepsilon}$; if so, one finds $S(n,p^{2};pc^{\prime}M)=-S(n,1;c^{\prime}M)$ only when $(c^{\prime},p)=1$, and zero otherwise. Our object (3.35) is degenerated into
@@
$\frac{1}{L}\sup_{C^{\prime}\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{\begin{subarray}{c}c^{\prime}\sim C^{\prime}\end{subarray}}\frac{S(n,1;c^{\prime}M)}{c^{\prime}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right).$ (3.37)
It thus suffices to deal with (3.35) with $(c,\ell)=1$ in what follows; by a similar argument the possibly-existing quantity (3.37) will provide a relative less significant contribution. Now, using (2.15), one reduces (3.35) essentially to
@@
Here, we have applied the estimate involving Ramanujan sum for each modulus $c$ that $S(n,0;c)=\sum_{ab=c}\mu(a)\sum_{\beta\bmod c}e{\left({\beta n}/{c}\right)}$. One should notice that the additive twist does not collude with the level anymore, namely, there holds that $(cM,q)=1$; this, however, is guaranteed by the second one of pre-conditions in (3.25) which confines that $c<q$ [^4] Indeed, on account of the final restrictions for the parameter $L$ (see (3.54) below), the pre-condition, here, seems dispensable (as far as the establishment of effective ranges for subconvexity is concerned).. By the asymptotic formula Lemma 2.2 for the integral transform $\Omega_{\pm}$, we are further reduced into the following superiorem problem that

### v4

@@ 2.8. Bessel functions
see, for example, [15, Proposition 8]. For any $x\in\mathbb{R}$, one might also write
@@
for any $j\geq 0$; see, for example, [48, Sections 16.12, 17.5 & 16.3]. Now, for any $h$, a smooth bump function supported on $[1/2,5/2]$ with bounded derivatives, we define
[…]
We thus have (see [24, Corollary 8.2]):
@@
[…]
@@
[…]
@@
[…]
@@
Remark 2.10.
Observe that the denominator $2\pi$ in both exponential factors of (2.18) and (2.19) usually looks inconspicuous in practical use. However, it plays a miraculous rôle in this paper, and would make it possible for us to apply the square-root estimate in Lemma 2.3. Otherwise, one cannot obtain the Lindelöf average bound in (1.6) whatsoever; see (4.7) and (4.11) below for details.
@@
Let $q$ be a prime and $M$ is also, satisfying that $(q,M)=1$. For any $f\in\mathcal{B}^{\ast}_{k}(M)$ and $F$ being a normalized $\text{GL}(3)$ Hecke-Maaß form of level $q^{2}$ with trivial nebentypus, we will consider the first moment of the central value $L(1/2,F\otimes f)$ in the levels aspects. Recall that $\mathcal{Q}_{F,f}\asymp M^{3}q^{4}$. The standard functional equation method (see, for example, [23, Theorem 5.3]) implies that the central value $L(1/2,F\otimes f)$ hinges on the following two quantities
$\begin{split}&\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}U{\left(\frac{n}{q^{2}M^{3/2}}\right)}\quad\text{and}\quad\varepsilon(F\otimes f)\sum_{n\geq 1}\frac{A_{\widetilde{F}}(n,1)\lambda_{f}(n)}{\sqrt{n}}V{\left(\frac{n}{q^{2}M^{3/2}}\right)},\end{split}$
@@
$L\geq\frac{q^{1+\varepsilon}}{{M^{1/4}}},\quad L\geq\frac{{M}^{1/4}}{{q^{1-\varepsilon}}}$ (3.2)
for any $\varepsilon>0$. We now define a set of modulo
$\mathscr{P}_{L}=\{p:p\text{ is prime, }L\leq p\leq 2L,(p,qM)=1\}.$ (3.3)
Let $f_{0}\in\mathcal{B}^{\ast}_{k}(M)$. By equipping with the amplifier (as in [6, Section 12.2])
$\mathfrak{A}_{f}=\sum_{j=1}^{2}\left|\sum_{\ell\in\mathscr{P}_{L}}\lambda_{f}(\ell^{j})x(\ell^{j})\right|^{2},$ (3.4)
with $x(n)={\text{sgn}}(\lambda_{f_{0}}(n))$, one attempts to estimate an amplified first moment
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\mathfrak{A}_{f}L(1/2,F\otimes f).$
It is clear that, by the relation $\lambda^{2}_{f_{0}}(p)-\lambda_{f_{0}}(p^{2})=1$, one sees that
$\mathfrak{A}_{f_{0}}\gg\frac{1}{2}\left(\sum_{\ell\in\mathscr{P}_{L}}|\lambda_{f_{0}}(\ell)|+|\lambda_{f_{0}}(\ell^{2})|^{2}\right)^{2}\gg\frac{L^{2}}{\log L}.$ (3.5)
If one opens the square and applies multiplicativity of Hecke eigenvalues, finally our main focus is on
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f),$ (3.6)
in view of the fact that
$\mathfrak{A}_{f}=\sum_{\ell\in\mathscr{P}_{L}}\left(x^{2}(\ell)+x^{2}(\ell^{2})\right)$ $+\sum_{\ell_{1},\ell_{2}\in\mathscr{P}_{L}}x(\ell^{2}_{1})x(\ell^{2}_{2})\lambda_{f}(\ell^{2}_{1}\ell^{2}_{2})$
$+\sum_{\ell_{1},\ell_{2}\in\mathscr{P}_{L}}\left(x(\ell_{1})x(\ell_{2})+x(\ell^{2}_{1})x(\ell^{2}_{2})\mathbf{1}_{\ell_{1}=\ell_{2}}\right)\lambda_{f}(\ell_{1}\ell_{2}).$ (3.7)
Notice that, here and thereafter,
={1,$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\},$ (3.8)
with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, is either 1 or a square-free integer or a square satisfying the coprimalities with the parameters $q$ and $M$, upon recalling (3.1).
@@
with $\mathcal{Q}^{(1-\delta)/2}_{F,f}\leq\mathcal{X}\leq\mathcal{Q}^{(1+\varepsilon)/2}_{F,f}$ for any small constants $\delta,\varepsilon>0$. Here, $U$ is a smooth bump function supported $[1/2,5/2]$, identical to one for $x\in[1,2]$, with derivatives satisfying that $x^{j}U^{(j)}\left({x}/{\mathcal{X}}\right)\ll_{j}1$ for any $j\geq 0$.
By the Cauchy-Schwarz inequality, together with the Rankin-Selberg bounds for the second moments of Fourier coefficients, the trivial estimate for the double sum in (3.9) is $\ll_{\varepsilon}\mathcal{X}^{1/2+\varepsilon}$. This is far away from the bound $\ll\mathcal{X}^{1/2+\varepsilon}/M$ leading to the convexity. The task is thus reduced to estimating (3.9) further. Our strategy next is the Petersson trace formula. To set up for our application of this, we trivially write the expression (3.9) above as
@@
It is straightforward to verify that, by the fact $|\lambda_{f}(M)|=M^{-1/2}$, the second term is bounded by $\ll_{\varepsilon}\sqrt{\mathcal{X}}/M^{3/2}$. For $(n\ell,M)=1$, notice that, by Lemma 2.6 one has the following relation
@@
Noticing that we are taking $k$ sufficiently large throughout the paper, one finds that the sum over $c$ can be truncated at $c\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$; the contribution in the complementary range is negligibly small (i.e., $O_{A}(\mathcal{X}^{-A})$ for any sufficiently large $A$), with the help of the bound (2.14). It is instructive to see that this sum of $c$ is not vacuous by the second pre-condition in (3.2); otherwise, one cannot succeed in obtaining hybrid subconvexity). In particular, whenever $\ell=1$, it requires that $q\geq M^{1/4+\varepsilon}$ (so that $\sqrt{\mathcal{X}}\gg M^{1+\varepsilon}$ for any $\varepsilon>0$), otherwise the display (3.14) is negligibly small; we will henceforth assume this pre-condition once we are considering the case of $\ell=1$ in the following paragraphs.
By using the dyadic subdivision, we further attempt to estimate the multiple sum for (3.14) as follows
@@
where ${U}_{\ell}\left(y\right)=U(y)J_{k-1}\left({4\pi\sqrt{y\mathcal{X}\ell}}/{CM}\right)$ for any $y\in\mathbb{R}^{+}$. According to the properties of the $J$-Bessel functions in §2.1, it follows that
@@
Next, we prepare to apply the Voronoĭ summation formula, Lemma 2.1. Before that, one needs to detect the coprimality between $\ell$ and the modulus of the twist $cM$. Recall that ={1,$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\}$ given as in (3.8).
Case I: $\ell=\ell_{1}\ell_{2}$. To begin with, let us consider what happens whenever $\ell=\ell_{1}\ell_{2}$. It is verifiable that the only counterexample with $(c,\ell)\neq 1$ occurs, provided that $q>M^{1/4+\varepsilon}$, and in which case, $c=\ell_{1}c^{\prime}$ or $\ell_{2}c^{\prime}$ for some $c^{\prime}\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}$, with $(c^{\prime},\ell_{1}\ell_{2}M)=1$ by first one of the restrictions in (3.2). If so, one finds that $S(n,\ell_{1}\ell_{2};\ell_{1}c^{\prime}M)=-S(n,\ell_{2}\overline{\ell_{1}};c^{\prime}M)$, provided that $(n,\ell_{1})=1$, and $S(n,\ell_{1}\ell_{2};\ell_{1}c^{\prime}M)=(\ell_{1}-1)S(n/\ell_{1},\ell_{2};c^{\prime}M)$ otherwise, where $\ell_{1}|n$. Our object (3.15) is essentially degenerated into two terms
@@
$\frac{1}{L}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\prime}\sim C}\frac{S(n,\ell_{2}\overline{\ell_{1}};c^{\prime}M)}{c^{\prime}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.17)
$\frac{A_{F}(\ell_{1},1)}{\sqrt{L}}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\prime}\sim C}\frac{S(n,\ell_{2};c^{\prime}M)}{c^{\prime}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}/L}\right);$ (3.18)
whilst, the expression of (3.15) we are concerned about is exactly
$\sup_{C\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c\sim C}\frac{S(n,\ell_{1}\ell_{2};cM)}{cM}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right).$ (3.19)
It suffices to evaluate (3.19) with $(c,\ell_{1}\ell_{2})=1$, observing that the expression (3.19) bears some resemblances to that in (3.17) and (3.18). By a similar argument, the two possibly-existing quantities (3.17) and (3.18) will provide a relative less significant contribution. For the sake of having a more clear understanding of the two quantities above for readers, we re-affirm that these two terms will not come into being, whenever $q<M^{1/4}$.
Case II: $\ell=\ell^{2}_{1}\ell^{2}_{2}$. Now, we turn to examining that the case of $\ell$ being a square in (3.15). Observe that, the (possibly occurred) scenarios where $(c,\ell)\neq 1$ are as follows:
  • (a)
    $\ell=\ell^{2}_{1}\ell^{2}_{2},c=c^{\ast}\ell_{1}$ or $c^{\ast}\ell_{2}$ for some $c^{\ast}\geq 1$, such that $(c^{\ast},\ell_{1}\ell_{2}M)=1$ and $c^{\ast}\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}$;
  • (b)
    $\ell=\ell^{2}_{1}\ell^{2}_{2},q>M^{1/4+\varepsilon}$, and $c=\hat{c}\ell_{1}\ell_{2}$ for some $\hat{c}\geq 1$, satisfying that $(\hat{c},\ell_{1}\ell_{2}M)=1$ and $\hat{c}\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}$.
Notice that, here, $\ell^{2}_{1}\ell_{1}^{2}|c$ and $\ell^{2}\ell^{2}_{2}|c$ can not occur, in view of the first constraint in (3.2). By a similar argument as before, one finds that, in those two cases, (a) and (b), the sum in (3.15) is thus degenerated into
$\frac{1}{L}$ $\sup_{C\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\ast}\sim C}\frac{S(n,\ell^{2}_{2};c^{\ast}M)}{c^{\ast}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.20)
$\frac{A_{F}(\ell_{1},1)}{\sqrt{L}}$ $\sup_{C\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\ast}\sim C}\frac{S(n,\ell_{1}\ell^{2}_{2};c^{\ast}M)}{\hat{c}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}/L}\right),$ (3.21)
and
$\frac{1}{L^{2}}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{\hat{c}\sim C}\frac{S(n,1;\hat{c}M)}{\hat{c}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.22)
$\frac{A_{F}(\ell_{1}\ell_{2},1)}{L}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{\hat{c}\sim C}\frac{S(n,\ell_{1}\ell_{2};\hat{c}M)}{\hat{c}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}/L^{2}}\right),$ (3.23)
respectively. Recall that, in Case II, the sum (3.15) is explicitly of the following form
$\sup_{C\ll\sqrt{\mathcal{X}}L^{2}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c\sim C}\frac{S(n,\ell^{2}_{1}\ell^{2}_{2};cM)}{cM}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right).$ (3.24)
On account of the similarities between the expression in (3.24) and that in (3.20)–(3.23), it suffices for us to handle the sum (3.24) with $(c,\ell_{1}\ell_{2})=1$, akin to Case I. In a similar vein, one can deal with the four sub-sums (3.20)–(3.23) in the degenerate case, and the contributions from them are dominated by that of (3.24). Upon combining with this and the argument in Case I as above, our goal eventually is boiled down to estimating (3.15) with $(c,\ell)=1$ in the following stage.
Having seen that the $n$-sum of (3.15) is already in a shape for an application of the the Voronoĭ, upon using (2.9), one might reduce (3.15) essentially to
@@
Here, we have applied the estimate involving Ramanujan sum for each modulus $c$ that $S(n,0;c)=\sum_{ab=c}\mu(a)\sum_{\beta\bmod c}e{\left({\beta n}/{c}\right)}$. One should notice that the additive twist does not collude with the level anymore, namely, there holds that $(cM,q)=1$; this, however, is guaranteed by the pre-condition in (3.1), which confines that $c<q$ [^3] Indeed, on account of the final restrictions for the parameter $L$ (see (3.46) below), the pre-condition, here, seems dispensable (as far as the establishment of effective ranges for subconvexity is concerned).. By the asymptotic formula, Lemma 2.2, for the integral transform $\Omega_{\pm}$, we are further reduced to the following superiorem problem that